An inductor of 250 mH has a reactance of 1450 Ω. What is the frequency of the current in the inductor?


1
Expert's answer
2021-06-29T08:43:59-0400

"X_L=2\pi f L"

"1450=2\pi \times f\times 250/1000=" "2\pi \times f\times 0.25"

"f=\frac{1450}{2\pi \times 0.25}= 923.0987 Hz"